# Oxyfluoride glasses obtained through incorporation of CaF2 into photovoltaic cover glass melts

### v2

@@ 2 Materials and methods
Reagents were mixed with CG in a platinum crucible and heated in a furnace, following a heating ramp consisting of 40 minutes at 200∘C, 60 minutes at 500∘C, 90 minutes at 800∘C, and 120 minutes at 1200∘C. After this stage, the molten glass was poured into a stainless steel mold and rapidly transferred to a second furnace maintained at 480∘C, where it was kept for approximately 12 hours to relieve internal stresses. Upon cooling to room temperature, the samples were removed from the mold. Figure 1 shows a photo of the obtained samples, which were cut, polished, or milled for the analyses.
The analytical techniques employed in this study encompass compositional, thermal, structural, and optical analyses. X-ray fluorescence (XRF) was used to determine the chemical composition of the samples. The analysis was performed on fused beads, and loss-on-ignition analysis was carried out to account for volatile content up to 1000∘C (Panalytical - Axios). Differential thermal analysis (DTA) was performed in the temperature range of 50∘C to 900∘C, using a heating rate of 10∘C/min under a synthetic air atmosphere with a flow rate of 50 mL/min (TA Instruments – SDT Q600). The density of the samples was obtained using Archimedes’ principle with distilled water at room temperature. Mass measurements were performed with an analytical balance with a precision of 0.1 mg (Kern ABT - 120-5DM).
Structural characterization was performed by X-ray powder diffraction (XRPD) on powdered samples. The analysis used CuK radiation (= 1.5418 Å) at 40 kV and 15 mA, with a scanning rate of 5∘/min, step size of 0.02∘, and a range of 5 to 80∘ (Rigaku - Miniflex 600). Also, in situ XRPD analysis was performed on powdered material to evaluate the temperature-dependent structural evolution. The measurements were conducted under air in the Bragg–Brentano geometry, using a 25∘C step size from 500 to 950∘C, over a range of 15 to 60∘, with a 0.016∘C step size and a counting time of 0.5 s/step. The powdered samples were placed in a platinum-lined corundum sample holder (Anton Paar - HTK1200N furnace with LynxEye detector).
@@ 3 Results and discussion
Assuming the CG composition obtained from the XRF analysis and the weights shown in Table 1, we can calculate the stoichiometric relations for CgCAF samples, which are very similar to those obtained experimentally by XRF. In this way, though fluorine is well known to be volatile, these results corroborate the assumption that no significant mass was lost during the melting process. As previously stated, we also aimed to keep the ratio SiO2/Na2O constant, and this XRF analysis confirms that this ratio is 3.5 for all samples.
Figure 2 presents the XRPD patterns for all glass samples analyzed in this study, where the dotted lines represent Lorentzian fits applied to the experimental data to support the subsequent discussion.
The absence of sharp peaks in all patterns confirms the amorphous nature of the samples, which is consistent with a glassy structural framework. These results agree with previously reported diffractograms for glasses and glass-ceramics of similar compositions [13, 21, 22]. Notably, the diffraction halos reveal the presence of two overlapping broad bands. In CG, the predominant band is centered at approximately 22∘, while for samples with increasing CaF2 content, the dominant feature progressively shifts towards 30∘. This behavior is likely associated with structural rearrangements driven by the incorporation of fluoride ions. As reported by Iwamoto and Makino [23], fluoride ions in calcium fluorosilicate systems can engage in Si-F or Ca-F bonding. At lower CaF2 concentrations, the formation of Si-F bonds is favored, whereas at higher concentrations, Ca-F interactions become more prominent. This transition is accompanied by a change in the fluoride coordination environment, from four-fold coordination, characteristic of a CaF2-type quasi-lattice, to six-fold coordination typical of a NaCl-type quasi-lattice. Notably, the NaCl-type structure possesses smaller average lattice parameters than the CaF2-type configuration. According to Bragg’s law, a reduction in the lattice spacing results in a shift of diffraction features towards higher 2 angles. Nonetheless, confirmation of this hypothesis requires more comprehensive structural analyses, including advanced modeling approaches.
@@
Figure 4 presents the thermograms of all samples, including CG, where the key events Tg (glass transition), Tx (onset of crystallization), and Tp (crystallization peak) are also indicated. However, in these experiments, Tx could not be identified for the CG, CgCAF05, and CgCAF12 samples.
As can be seen, in all CgCAF samples, it is possible to identify two crystallization peaks, a result widely observed in the literature on silicates containing CaF2 [13, 16], as well as a significant reduction in the glass transition temperature (50-90∘C). A summary of the characteristic temperatures identified for all samples is shown in Table 3.
@@
Consistent with the DTA measurements, a crystalline phase appears at 600∘C. However, we cannot identify this phase using the ICDD database. Additional work will be necessary to elucidate this point. The relative intensity of the diffraction peaks associated with the unidentified phase increases progressively up to 800∘C, indicating its growth, followed by a gradual decline in intensity at higher temperatures. Nonetheless, traces of this phase remain detectable at 950∘C. At 650∘C the crystallization of tricalcium silicate C3S (Ca3SiO5) is observed; however, this phase vanishes upon further heating and is no longer detectable at 800∘C. The Combeite phase (Na4Ca4Si6O18) crystallizes at around 675∘C, and its proportion increases until 950∘C at which point it becomes the dominant crystalline phase at high temperature. Following the experiment, the XRPD pattern collected at room temperature, after thermal cycling, shows the presence of Combeite phase, along with trace amounts of Akermanite. Some of these phases have been linked to intriguing features, such as antimicrobial activity [24]. This indicates that CG-derived materials may also result in glass-ceramics with unique properties, which we will explore in future work.

### v3

@@ Abstract
[IFPR]organization=Instituto Federal do Paraná,city=Barracão, state=Paraná, country=Brasil
@@ 2 Materials and methods
Reagents were mixed with CG in a platinum crucible and heated in a furnace, using a 10 ∘C/min heating rate, with holding steps of 40 minutes at 200∘C, 60 minutes at 500∘C, 90 minutes at 800∘C, and 120 minutes at 1200∘C. This thermal profile was chosen to ensure a homogeneous fusion of the mixture and to avoid foaming or spillage during melting. After this stage, the molten glass was poured into a stainless steel mold and rapidly transferred to a second furnace maintained at 480∘C, where it was kept for approximately 12 hours to relieve internal stresses. Upon cooling to room temperature, the samples were removed from the mold. Figure 1 shows a photo of the obtained samples, which were cut, polished, or milled for the analyses.
The analytical techniques employed in this study encompass compositional, thermal, structural, and optical analyses. X-ray fluorescence (XRF) was used to determine the chemical composition of the samples. The analysis was performed on fused beads at approximately 1200∘C (Panalytical – Axios). The beads were prepared from powdered material, previously ground to pass through a 325-mesh sieve, using lithium tetraborate as the flux. Additionally, loss-on-ignition analysis was carried out to account for volatile content up to 1000∘C. Differential thermal analysis (DTA) was performed in the temperature range of 50∘C to 900∘C, using a heating rate of 10∘C/min under a synthetic air atmosphere with a flow rate of 50 mL/min (TA Instruments – SDT Q600). The density of the samples was obtained using Archimedes’ principle with distilled water at room temperature. Mass measurements were performed with an analytical balance with a precision of 0.1 mg (Kern ABT - 120-5DM).
Structural characterization was performed by X-ray powder diffraction (XRPD) on powdered samples. The analysis used CuK radiation (= 1.5418 Å) at 40 kV and 15 mA, with a scanning rate of 5∘/min, step size of 0.02∘, and a range of 5 to 80∘ (Rigaku - Miniflex 600). Also, in situ XRPD analysis was performed on powdered material to evaluate the temperature-dependent structural evolution. The measurements were conducted under air in the Bragg–Brentano – geometry, using Ni-filtered CuK radiation ( = 1.5418Å), over a 2 range of 15 to 60∘, with a step size of 0.016∘ and a counting time of 0.5 s/step. A temperature step of 25∘C was applied from 500 to 950∘C. Data were collected on a Bruker-AXS D8 Advance diffractometer (Karlsruhe, Germany) equipped with an Anton Paar oven chamber (model HTK 1200N, Graz, Austria), which allows measurements up to 1473 K. The powdered samples were placed in a platinum-lined corundum sample holder, and the diffracted intensities were recorded using a LynxEye detector.
@@ 3 Results and discussion
Assuming the CG composition obtained from the XRF analysis and the weights shown in Table 1, we can calculate the stoichiometric relations for CgCAF samples, which are very similar to those obtained experimentally by XRF. In this way, though fluorine is well known to be volatile, these results corroborate the assumption that no significant mass was lost during the melting process. As previously stated, we also aimed to keep the ratio SiO2/Na2O constant, and the XRF analysis confirms this it lies in the range of 3.5 - 3.7 for the samples.
Figure 2 presents the XRPD patterns for all glass samples. The absence of sharp peaks in all patterns confirms the amorphous nature of the samples, consistent with a glassy structural framework. These results are in agreement with previously reported diffractograms for glasses and glass-ceramics of similar compositions [13, 21, 22].
Notably, the diffraction halos exhibit two overlapping broad bands, which can be more clearly distinguished by fitting the patterns with two Lorentzian curves, as shown in 2, represented by the filled green and blue curves corresponding to bands centered at aprproximately 22∘ and 30∘, respectivelly. The area ratio between the green and blue bands decreases from 1.16 for the CG sample to 0.69 to the CgCAF20 sample. In the CG sample, the most intense band is centered around 22∘, whereas in samples with increasing CaF2 content, the band at 30∘ becomes more prominent. This behavior is likely associated with structural rearrangements driven by the incorporation of fluoride ions. As reported by Iwamoto and Makino [23], fluoride ions in calcium fluorosilicate systems can engage in Si-F or Ca-F bonding. At lower CaF2 concentrations, the formation of Si-F bonds is favored, whereas at higher concentrations, Ca-F interactions become more prominent. This transition is accompanied by a change in the fluoride coordination environment, from four-fold coordination, characteristic of a CaF2-type quasi-lattice, to six-fold coordination typical of a NaCl-type quasi-lattice. Notably, the NaCl-type structure possesses smaller average lattice parameters than the CaF2-type configuration. According to Bragg’s law, a reduction in the lattice spacing results in a shift of diffraction features towards higher 2 angles. Nonetheless, confirmation of this hypothesis requires more comprehensive structural analyses, including advanced modeling approaches.
@@
Figure 4 presents the thermograms of all samples, including CG, where the key events Tg (glass transition), Tx (onset of crystallization), and Tp (crystallization peak) are also indicated. A summary of the characteristic temperatures identified for all samples is shown in Table 3. However, in these experiments, Tx could not be identified for the CG, CgCAF05, and CgCAF12 samples.
As can be seen, in all CgCAF samples, it is possible to identify two crystallization peaks, a result widely observed in the literature on silicates containing CaF2 [13, 16], as well as a significant reduction in the glass transition temperature (50-90∘C).
@@
Consistent with DTA measurements, a crystalline phase appears at 600∘C. Nevertheless, we cannot identify this phase using the ICDD PDF5 database. Additional work will be necessary to elucidate this point. The proportion of this unknown phase increases until 800∘C and after decreases but is present in small amounts at 950∘C. At 650∘C we observe the crystallization of C3S (Ca3SiO5 JCPDS file 00-017-0445), which disappears at 800∘C. The Combeite phase (Na4Ca4Si6O18 JCPDS file 04-008-0810) crystallizes around 700∘C, and its proportion increases until 950∘C to be finally the main phase at high temperature. After this experiment, the PXRD diagram collected at room temperature shows the presence of Combeite and a minimal amount of Akermanite (JCPDS file 01-079-2425). Some of these phases have been linked to intriguing features, such as antimicrobial activity [24]. This indicates that CG-derived materials may also result in glass-ceramics with unique properties, which we will explore in future work.